OpenStudy (punderson):

The function f(x) = -x2 - 9x + 10 shows the relationship between the vertical distance of a diver from a pool's surface f(x), in feet, and the horizontal distance x, in feet, of a diver from the diving board. What is a zero of f(x), and what does it represent? (4 points) x = 1; the diver jumps in the pool at 1 foot per second x = 10; the diver jumps in the pool at 10 feet per second x = 1; the diver hits the water 1 foot away horizontally from the board x = 10; the diver hits the water 10 feet away horizontally from the board

OpenStudy (punderson):

@Error1603

OpenStudy (punderson):

What are the exact solutions of x2 - 5x - 7 = 0? (4 points) x = x = x = x =

OpenStudy (punderson):

hold on ill fix that^

OpenStudy (error1603):

x = 10; the diver hits the water 10 feet away horizontally from the board

OpenStudy (error1603):

Use \[-b \pm b2-4ac-------√2a\]

OpenStudy (error1603):

\[-b \pm b2 -4ac \sqrt{2a}\]

OpenStudy (error1603):

Last one Is D

OpenStudy (punderson):

Thank you, i might have 2 more, if you dont mind

OpenStudy (error1603):

ok

OpenStudy (punderson):

23. If y = v - 0.5gt2, which equation is solved for t? (3 points) 24. An equation is shown below: km2 + cp = 2 Which equation is solved for m? (3 points)

OpenStudy (punderson):

top one is for 23